# Current address: ]College of William and Mary, Williamsburg, Virginia 23187-8795

### v2

@@ VII.6.2 Isospin Study of $\overline{\gamma}_{0}(Q^{2})$
Our results on the isovector and isoscalar combination of $\overline{\gamma}_{0}(Q^{2})$ from EG4 are shown in Fig. 26. We also provide results based on EG4 $p$ data and E97110 $n$(3He) data [104]. The total experimental systematic uncertainties on the proton and deuteron/neutron(3He) are combined quadratically, as are the statistical uncertainties. The low-$x$ uncertainties are added linearly. The results shown in Fig. 26 agree reasonably well with those from the previous experiments EG1b and EG1b/E94-010 [43]. However, unlike for $\overline{\Gamma}_{1}^{p-n}$, results from the EG4-only and EG4/E97110 approaches agree only at the higher $Q^{2}$ points. On the other hand, if we combine linearly the total systematic uncertainties of the two experiments, the tension at low $Q^{2}$ vanishes.
For $\overline{\gamma}_{0}^{p-n}(Q^{2})$, both the EG4 and EG4/E97110 results indicate that it stays positive in the measured $Q^{2}$ range, conversely to the $\chi$EFT and MAID predictions. In the case of $\overline{\gamma}_{0}(Q^{2})^{p+n}$, both the EG4 and EG4/E97110 combinations agree with Bernard et al. in the lower $Q^{2}$ range, but disagree with the Alarcón et al. prediction.
Figure 26: [Color online] EG4 results on the isovector (top) and isoscalar (bottom) combinations of the generalized longitudinal spin polarizability $\overline{\gamma}_{0}(Q^{2})$. The EG4-only data are shown by the red circles, compared with the earlier EG1b-only (black crossbars), the combination of EG4 $p$ and E97110 $n$(3He) data (blue triangles), and the combination of EG1b $p$ and E94010 $n$(3He) data (green triangles). For $\overline{\gamma}_{0}^{p-n}(Q^{2})$, the size of the systematic uncertainty varies from $0.026$ to $0.522$, while the relative size of the systematic to statistical uncertainty varies from $(58-216)\%$. For $\overline{\gamma}_{0}^{p+n}(Q^{2})$, the size of the systematic uncertainty varies from $0.013$ to $0.367$, while the relative size of the systematic to statistical uncertainty varies from $(36-93)\%$. The results are compared with $\chi$EFT predictions from Alarcón et al. [60] and Bernard et al., as well as the MAID model. The prediction by Bernard et al. [59] does not provide uncertainty bands for the proton-neutron difference; for illustration, we show a band whose width is the larger of the proton and the neutron uncertainties.

### v3

@@ VII.6.2 Isospin Study of $\overline{\gamma}_{0}(Q^{2})$
Our results on the isovector and isoscalar combination of $\overline{\gamma}_{0}(Q^{2})$ from EG4 are shown in Fig. 26. We also provide results based on EG4 $p$ data and E97-110 $n$(3He) data [104]. The total experimental systematic uncertainties on the proton and deuteron/neutron(3He) are combined quadratically, as are the statistical uncertainties. The low-$x$ uncertainties are added linearly. The results shown in Fig. 26 agree reasonably well with those from the previous experiments EG1b and EG1b/E94-010 [43]. However, unlike for $\overline{\Gamma}_{1}^{p-n}$, results from the EG4-only and EG4/E97-110 approaches agree only at the higher $Q^{2}$ points. On the other hand, if we combine linearly the total systematic uncertainties of the two experiments, the tension at low $Q^{2}$ vanishes.
For $\overline{\gamma}_{0}^{p-n}(Q^{2})$, both the EG4 and EG4/E97-110 results indicate that it stays positive in the measured $Q^{2}$ range, conversely to the $\chi$EFT and MAID predictions. In the case of $\overline{\gamma}_{0}(Q^{2})^{p+n}$, both the EG4 and EG4/E97-110 combinations agree with Bernard et al. in the lower $Q^{2}$ range, but disagree with the Alarcón et al. prediction.
Figure 26: [Color online] EG4 results on the isovector (top) and isoscalar (bottom) combinations of the generalized longitudinal spin polarizability $\overline{\gamma}_{0}(Q^{2})$. The EG4-only data are shown by the red circles, compared with the earlier EG1b-only (black crossbars), the combination of EG4 $p$ and E97-110 $n$(3He) data (blue triangles), and the combination of EG1b $p$ and E94-010 $n$(3He) data (green triangles). For $\overline{\gamma}_{0}^{p-n}(Q^{2})$, the size of the systematic uncertainty varies from $0.026$ to $0.522$ (×$\times 10^{-4}$ fm4), while the relative size of the systematic to statistical uncertainty varies from $(58-216)\%$. For $\overline{\gamma}_{0}^{p+n}(Q^{2})$, the size of the systematic uncertainty varies from $0.013$ to $0.367$ (×$\times 10^{-4}$ fm4), while the relative size of the systematic to statistical uncertainty varies from $(36-93)\%$. The results are compared with $\chi$EFT predictions from Alarcón et al. [60] and Bernard et al., as well as the MAID model. The prediction by Bernard et al. [59] does not provide uncertainty bands for the proton-neutron difference; for illustration, we show a band whose width is the larger of the proton and the neutron uncertainties.